One side of a seesaw carries a 21kg mass four meters from the fulcrum and a 25.5 kg mass two meters from the fulcrum.

Where do you need to place a 35 kg mass on the side opposite of the first two masses to balance the seesaw?
Answer #1

let the 35 kg mass be placed at a distance x from the fulcrum

taking moment about the fulcrum we have

m1x1 + m2x2 = m3x3

21*4 + 25.5*2 = 35*x

x = 3.86 m

so the 35 kg mass is placed 3.86 m to the right